College students average 9 hours of sleep per night with a standard deviation of 45 minutes. If the amount of sleep is normally distributed, what proportion of college students sleep for more than 10.1 hours?
Answer #1

Normal distribution: P(X < A) = P(Z < (A - mean)/standard deviation)

Mean = 9 hours

Standard deviation = 45 minutes = 0.75 hours

P(a student sleeps for more than 10.1 hours), P(X > 10.1) = 1 - P(X < 10.1)

= 1 - P(Z < (10.1 - 9)/0.75)

= 1 - P(Z < 1.47)

= 1 - 0.9292

= 0.0708